OpenStudy (anonymous):

Need answer to this question on Capacitor so that i can compare my solutions to yours. Voltage across a capacitor of 2 F is given by V(t) = 5+ SinSquare2t. a. Find an expression for i(t). b. Find the power delivered to the capacitor at 10s. c. Find the maximum energy stored in the capacitor @Physics

OpenStudy (anonymous):

Is it \((v(t)=5+\sin(\sqrt{2}t)\)? Is t under the square root or not?

OpenStudy (anonymous):

I believe it is 5*sin instead of (5+sin) am i right? Let's say: V(t) = 5* Sin(Square(2)t) = 5* Sin(Square(2)t + 0) = 5 < 0 , in phasor form So the current [ i(t) ] must be lagged 90 degrees or pi/2 rad from the voltage. Unless you have a trifasic system.. then you have to add 120 degrees Therefore: a) i(t) = 5 < 0 + 90 ,for example. i(t) = 5<90 b) V(10) = 5* Sin(Square(2)*10) = 1.22 V c) how can i know if i dont have the capacitor properties?